# Elliptic asymptotic representation of the fifth Painlevé transcendents (corrected version)

### v8

@@ 5.1. Expressions in terms of elliptic integrals
$J^{\mu}_{2}-\hat{J}^{\mu}_{2}$ $={e^{i\phi}}t\int_{\mathbf{b}}\mu(\lambda(z))dz$
$=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\Bigl(\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}+\frac{2e^{-i\phi}\theta_{\infty}t^{-1}z}{w}\Bigr)dz+{e^{i\phi}}t^{-1}\int_{\mathbf{b}}\tilde{g}_{2}(t,z)dz$
$=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}dz-\frac{\theta_{\infty}\pi i}{2}+O(t^{-1}).$
Observing that $\sqrt{(a_{\phi}-z^{2})/(1-z^{2})}=(1/w)(a_{\phi}-z^{2})=-(w/z)^{\prime}+a_{\phi}/w-z^{-2}a_{\phi}/w,$ we have the following.
@@
$J^{\mu}_{2}-\hat{J}^{\mu}_{2}$ $=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}dz-\frac{\theta_{\infty}\pi i}{2}+O(t^{-1})$
$=\frac{e^{i\phi}}{4}a_{\phi}t\int_{\mathbf{b}}\Bigl(\frac{1}{w}-\frac{1}{z^{2}w}\Bigr)dz-\frac{\theta_{\infty}\pi i}{2}+O(t^{-1}).$
To calculate $J_{1,2},$ $\hat{J}_{2}$ it is necessary to know integrals of
@@
and that $\mu(z_{\pm})^{2}=b_{3}(z_{\pm})^{2}.$ Furthermore, by (5.3), $\|\mathrm{diag}T^{-1}T_{\lambda}\|\ll|z\mp 1|^{-1/2}$ near $z=\pm 1,$ $\mathrm{diag}T^{-1}T_{\lambda}$ has poles at $z_{\pm}\in\Pi_{-}$, and is holomorphic around $z_{\pm}\in\Pi_{+}.$ From (3.9) combined with (5.2) it follows that
@@
These combined with Proposition 5.2 and (4.4) yield
@@
$W_{0}(z)=\frac{e^{i\phi}}{4}a_{\phi}\Bigl(\frac{1}{w}-\frac{1}{z^{2}w}\Bigr),\quad W_{1}(z)=\frac{1}{4}\Bigl(\frac{z_{+}-z_{-}}{w}+\frac{w(z_{+})}{(z-z_{+})w}-\frac{w(z_{-})}{(z-z_{-})w}\Bigr).$
Then
$J_{2}-\hat{J}_{2}$ $=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}dz-\int_{\mathbf{b}}W_{1}(z)dz-\frac{\theta_{\infty}\pi i}{2}-2\pi ir_{\mathbf{b}}+O(t^{-1})$
$=\int_{\mathbf{b}}(tW_{0}(z)-W_{1}(z))dz-\frac{\theta_{\infty}\pi i}{2}-2\pi ir_{\mathbf{b}}+O(t^{-1}),$
@@ 5.2. Expressions in terms of the $\vartheta$-function
Then $z=\mathrm{\check{s}n}(u)$ satisfies $(dz/du)^{2}=w(z)^{2}=(1-z^{2})(a_{\phi}-z^{2}).$ Setting $z=\mathrm{\check{s}n}(u)$ we have, for a given $z_{0}$ on the elliptic curve $\Pi=\Pi_{+}\cup\Pi_{-}$,
@@
$\vartheta(z\pm 1,\tau)=\vartheta(z,\tau),\quad\vartheta(-z,\tau)=\vartheta(z,\tau),\quad\vartheta(z\pm\tau,\tau)=e^{-\pi i(\tau\pm 2z)}\vartheta(z,\tau).$
Let us write the fundamental periods of $\Pi$ as
$\omega_{\mathbf{a}}=\omega_{\mathbf{a}}(t)=\int_{\mathbf{a}}\frac{dz}{w},\quad\omega_{\mathbf{b}}=\omega_{\mathbf{b}}(t)=\int_{\mathbf{b}}\frac{dz}{w}.$
@@
$-\frac{1}{w(z_{0}^{(+)})}\Bigl(\frac{w^{\prime}(z_{0}^{(+)})}{2}-\frac{1}{\omega_{\mathbf{a}}}\Bigl({i\pi}+\frac{\vartheta^{\prime}}{\vartheta}(F(z^{(-)}_{0},z^{(+)}_{0})+\nu,\tau)\Bigr)\Bigr)\frac{dz}{w},$
@@
$\Bigl(\frac{z_{0}}{2}-\frac{w^{\prime}(z_{0})}{4}\Bigr)\omega_{\mathbf{a}}+\frac{1}{2}\frac{\vartheta^{\prime}}{\vartheta}(F(z^{(-)}_{0},z^{(+)}_{0})+\nu,\tau)+\frac{\pi i}{2}=\frac{\vartheta^{\prime}}{\vartheta}\Bigl(\frac{1}{2}F(z^{(-)}_{0},z^{(+)}_{0})-\frac{1}{4},\tau\Bigr)$
derived by comparing the residues of poles $z_{0}=\pm 1,$ $\pm a_{\phi}^{1/2}$ and $\infty^{+}$ $(\in\Pi_{+})$ on $\Pi$ (cf. [22, p.513], [28, (3.5)]), we obtain from (5.8) that
$\int_{\mathbf{a}}\frac{w(z_{0})dz}{(z-z_{0})w(z)}=-z_{0}\omega_{\mathbf{a}}+2\frac{\vartheta^{\prime}}{\vartheta}\Bigl(\frac{1}{2}F(z^{(-)}_{0},z^{(+)}_{0})-\frac{1}{4},\tau\Bigr).$
In what follows let us adopt the convention that the path of the integral $\omega_{\mathbf{a}}F(z^{(-)}_{0},z^{(+)}_{0})=\int^{z^{(+)}_{0}}_{z^{(-)}_{0}}w(z)^{-1}dz$ passes through $a_{\phi}^{1/2}$, i.e. the left end of the cut $a^{1/2}_{\phi},1]$. Then
@@
and hence $F(-z^{(-)}_{0},-z^{(+)}_{0})=F(z^{(-)}_{0},z^{(+)}_{0})+1.$ From these relations with $_{0}=z_{+}, zz_{-}$ such that $z_{+}+z_{-}=O(t^{-1})$ (cf. (5.5)) it follows that
@@ 5.3. Expression of $B_{\phi}(t)$
Recall that $a_{\phi}(t)=A_{\phi}+t^{-1}B_{\phi}(t),$ $B_{\phi}(t)=O(1)$ in the domain $S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})$, where $A_{\phi}$ is a unique solution of the Boutroux equation (2.3). The quantity $B_{\phi}(t)$ is written in terms of
@@
where $w(A_{\phi},z)=\sqrt{(A_{\phi}-z^{2})(1-z^{2})},$ and $\mathbf{a},$ $\mathbf{b}$ are two cycles on $\Pi^{*}$. Observing that
@@
In $S^{\prime}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})$, $B_{\phi}(t)$ is bounded, and
@@
in which $\hat{F}$ denotes $F$ corresponding to $\hat{\tau}=(-\omega_{\mathbf{a}})/\omega_{\mathbf{b}}.$ Since $\mathrm{Re\,}\int_{\mathbf{a},\mathbf{b}}W_{1}(z)dz$ are bounded in $S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1}),$ the Boutroux equations (2.3) with $A_{\phi}$ are equivalent to the boundedness of $\mathrm{Re\,}(e^{i\phi}\Omega_{\mathbf{a},\mathbf{b}}B_{\phi})$, namely, the boundedness of $B_{\phi}(t).$
@@
The derivation of the last equality is similar to that of Legendre’s relation [14], [43].
@@
By the boundedness of $B_{\phi}(t),$ $\omega_{\mathbf{a},\,\mathbf{b}}=\Omega_{\mathbf{a},\,\mathbf{b}}+O(t^{-1})$ for $\mathbf{a},$ $\mathbf{b}\subset\Pi\cap\Pi^{*}.$ From Proposition 5.5 and Corollary 5.4, it follows that
@@
Observe that the contour $z_{+}^{(-)},z_{+}^{(+)}]^{\sim}$ on $\Pi\cap\Pi^{*}$ may be decomposed into $2p_{+}(t)\mathbf{a}\cup 2q_{+}(t)\mathbf{b}\cup\mathbf{a}_{+}\cup\mathbf{a}_{-},$ where the length of $\mathbf{a}_{\pm}$ is $\ll 1$. Using (5.10) and Lemma 5.8, we have
@@
$\ll$ $|\mathcal{E}_{\mathbf{b}}j_{\mathbf{a}}-\mathcal{E}_{\mathbf{a}}j_{\mathbf{b}}|+O(t^{-1})=\Bigl|\frac{\partial}{\partial A_{\phi}}(\mathcal{E}_{\mathbf{a}}\Omega_{\mathbf{b}}-\mathcal{E}_{\mathbf{b}}\Omega_{\mathbf{a}})\Bigr|+O(t^{-1})\ll t^{-1}$
@@
Let $y(t)$ be a function satisfying (4.3), and let $(M^{0},M^{1})\in SL_{2}(\mathbb{C})^{2}$ be such that $m^{0}_{11}m^{1}_{11}m^{0}_{21}m^{1}_{12}\not=0.$ Suppose that $0<|\phi|<\pi/2.$
@@
By the same argument as in the derivation of the formula in Proposition 5.6 we have
@@
Justification The justification of (6.3) as a solution of (PV) is made along the argument in [28, pp. 105–106, pp. 120–121]. Let $\mathcal{M}=(\mathfrak{m}_{\phi},m^{0}_{21}m^{1}_{12})$ be a point such that $m^{0}_{11}m^{1}_{11}m^{0}_{21}m^{1}_{12}\not=0$ on the monodromy manifold for system (3.4). Relation (6.3) and Corollary 6.1 provide the leading terms $y_{\mathrm{as}}=y_{\mathrm{as}}(\mathcal{M},t)$ and $(B_{\phi})_{\mathrm{as}}=(B_{\phi})_{\mathrm{as}}(\mathcal{M},t)$ without $O(t^{-\delta})$. Viewing (4.3) we set $2y^{*}_{\mathrm{as}}=2y^{*}_{\mathrm{as}}(\mathcal{M},t)=\sqrt{\varphi(t,y_{\mathrm{as}},(B_{\phi})_{\mathrm{as}})}$ with
$\varphi(t,y,B)=$ $e^{2i\pi}y(4y+(1-A_{\phi})(y-1)^{2})$
@@
where the branch of the square root is chosen in such a way that the leading term of $y^{*}_{\mathrm{as}}$ is compatible with $(d/dt)y_{\mathrm{as}}$. Then $(y_{\mathrm{as}},y^{*}_{\mathrm{as}})=(y_{\mathrm{as}}(\mathcal{M},t),y^{*}_{\mathrm{as}}(\mathcal{M},t))$ satisfies (4.3) with $B_{\phi}(t)=(B_{\phi})_{\mathrm{as}}(\mathcal{M},t)$ in $\hat{S}(\phi,t_{\infty},\kappa_{0},\delta_{2})$, and the relation $\mathcal{M}=\mathcal{M}(t,y_{\mathrm{as}},y^{*}_{\mathrm{as}})$ [28, (2.28)], where $\mathcal{M}(t,y,y^{*})$ is a collection of explicit functions $y(t),$ $B_{\phi}(t)$ resulting from the WKB procedure. Here $\hat{S}(\phi,t_{\infty},\kappa_{0},\delta_{2})=\{t\,|\,\,\mathrm{Re\,}t>t_{\infty},|\mathrm{Im\,}t|<\kappa_{0}\}\setminus\bigcup_{\sigma\in Z_{\pm 1}\cup Z_{\infty}\cup Z_{0}}\{|t-e^{-i\phi}\sigma|\leq\delta_{2}\}$ with $Z_{\pm 1}=x_{0}+\Omega_{\mathbf{a}}(\mathbb{Z}+\tfrac{1}{2})+2\Omega_{\mathbf{b}}\mathbb{Z},$ $Z_{\infty}=x_{0}+\Omega_{\mathbf{a}}\mathbb{Z}+2\Omega_{\mathbf{b}}(\mathbb{Z}+\tfrac{1}{2}),$ $Z_{0}=\{e^{i\phi}t\,|\,\,y_{\mathrm{as}}(t)=0,\infty\}.$ The monodromy data $\mathcal{M}_{\mathrm{as}}(t)$ for system (3.4) with $\mathcal{B}(t,\lambda)$ containing $(y_{\mathrm{as}},y^{*}_{\mathrm{as}})$ is given by $\mathcal{M}_{\mathrm{as}}(t)=\mathcal{M}(t,y_{\mathrm{as}},y^{*}_{\mathrm{as}})+O(t^{-\delta})$ [28, (2.27)] as a result of the repeated WKB procedure. Thus, for $|t|\geq T_{0}$, we have $\|\mathcal{M}_{\mathrm{as}}(t)-\mathcal{M}\|\leq Ct^{-\delta}$ valid uniformly in a neighbourhood of $\mathcal{M}$, where $C$ and $T_{0}$ are independent of $\mathcal{M}$ [28, (2.26)]. Then the justification scheme [26] applies to our case (see also [12, Theorem 5.5]). This justification combined with the maximal modulus principle in each neighbourhood of $\sigma\in Z_{0}\cup Z_{\pm 1}$ leads us Theorem 2.1 in $S(\rho,t_{\infty},\kappa_{0},\delta_{0}).$

### v9

@@ 5.1. Expressions in terms of elliptic integrals
$J^{\mu}_{2}-\hat{J}^{\mu}_{2}$ $={e^{i\phi}}t\int_{\mathbf{b}}\mu(\lambda(z))dz+O(t^{-1/2})$
$=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\Bigl(\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}+\frac{2e^{-i\phi}\theta_{\infty}t^{-1}z}{w}\Bigr)dz+{e^{i\phi}}t^{-1}\int_{\mathbf{b}}\tilde{g}_{2}(t,z)dz+O(t^{-1/2})$
$=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}dz-\frac{\theta_{\infty}\pi i}{2}+O(t^{-1/2}).$
The relation $\sqrt{(a_{\phi}-z^{2})/(1-z^{2})}=(1/w)(a_{\phi}-z^{2})=-(w/z)^{\prime}+a_{\phi}/w-z^{-2}a_{\phi}/w$ yields the following.
@@
$J^{\mu}_{2}-\hat{J}^{\mu}_{2}$ $=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}dz-\frac{\theta_{\infty}\pi i}{2}+O(t^{-1/2})$
$=\frac{e^{i\phi}}{4}a_{\phi}t\int_{\mathbf{b}}\Bigl(\frac{1}{w}-\frac{1}{z^{2}w}\Bigr)dz-\frac{\theta_{\infty}\pi i}{2}+O(t^{-1/2}).$
To calculate $J_{1,2},$ $\hat{J}_{2}$ it is necessary to know the integrals of
@@
and that $\mu(z_{\pm})^{2}=b_{3}(z_{\pm})^{2}.$ By (5.3), $\|\mathrm{diag}T^{-1}T_{\lambda}\|\ll|z\mp 1|^{-1/2}$ near $z=\pm 1.$ Furthermore, $\mathrm{diag}T^{-1}T_{\lambda}$ has poles at $z_{\pm}\in\Pi_{-}$ and is holomorphic around $z_{\pm}\in\Pi_{+}.$ From (3.9) combined with (5.2) it follows that
@@
These combined with Proposition 5.2 and (4.4) yield the following.
@@
$W_{0}(z)=\frac{e^{i\phi}}{4}a_{\phi}\Bigl(\frac{1}{w}-\frac{1}{z^{2}w}\Bigr),\quad W_{1}(z)=\frac{1}{4}\Bigl(\frac{z_{+}-z_{-}}{w}+\frac{w(z_{+})}{(z-z_{+})w}-\frac{w(z_{-})}{(z-z_{-})w}\Bigr)$
with $w=w(z)=w(a_{\phi},z).$ Then
$J_{2}-\hat{J}_{2}$ $=\frac{e^{i\phi}}{4}t\int_{\mathbf{b}}\sqrt{\frac{a_{\phi}-z^{2}}{1-z^{2}}}dz-\int_{\mathbf{b}}W_{1}(z)dz-\frac{\theta_{\infty}\pi i}{2}-2\pi ir_{\mathbf{b}}+O(t^{-1/2})$
$=\int_{\mathbf{b}}(tW_{0}(z)-W_{1}(z))dz-\frac{\theta_{\infty}\pi i}{2}-2\pi ir_{\mathbf{b}}+O(t^{-1/2}),$
@@ 5.2. Expressions in terms of the $\vartheta$-function
Then $z=\mathrm{\check{s}n}(u)$ satisfies $(dz/du)^{2}=w(z)^{2}=(1-z^{2})(a_{\phi}-z^{2}).$ Setting $z=\mathrm{\check{s}n}(u)$ we have, for a given $z_{0}$ on the elliptic curve $\Pi_{a_{\phi}}=\Pi_{a_{\phi},+}\cup\Pi_{a_{\phi},-}$ for $w(z)=w(a_{\phi},z)$,
@@
$\vartheta(z\pm 1,\tau)=\vartheta(z,\tau),\quad\vartheta(-z,\tau)=\vartheta(z,\tau),\quad\vartheta(z\pm\tau,\tau)=e^{-\pi i(\tau\pm 2z)}\vartheta(z,\tau)$
[14, 43]. Let us write the fundamental periods of $\Pi_{a_{\phi}}$ as
$\omega_{\mathbf{a}}=\omega_{\mathbf{a}}(t)=\int_{\mathbf{a}}\frac{dz}{w(z)},\quad\omega_{\mathbf{b}}=\omega_{\mathbf{b}}(t)=\int_{\mathbf{b}}\frac{dz}{w(z)},\quad w(z)=w(a_{\phi},z).$
@@
$-\frac{1}{w(z_{0}^{(+)})}\Bigl(\frac{w^{\prime}(z_{0}^{(+)})}{2}-\frac{1}{\omega_{\mathbf{a}}}\Bigl({i\pi}+\frac{\vartheta^{\prime}}{\vartheta}(F(z^{(-)}_{0},z^{(+)}_{0})+\nu,\tau)\Bigr)\Bigr)\frac{dz}{w(z)},$
@@
$\int_{\mathbf{a}}\frac{w(z_{0})dz}{(z-z_{0})w(z)}=-z_{0}\omega_{\mathbf{a}}+2\frac{\vartheta^{\prime}}{\vartheta}\Bigl(\frac{1}{2}F(z^{(-)}_{0},z^{(+)}_{0})-\frac{1}{4},\tau\Bigr),$
combining (5.8) with the relation
$\Bigl(\frac{z_{0}}{2}-\frac{w^{\prime}(z_{0})}{4}\Bigr)\omega_{\mathbf{a}}+\frac{1}{2}\frac{\vartheta^{\prime}}{\vartheta}(F(z^{(-)}_{0},z^{(+)}_{0})+\nu,\tau)+\frac{\pi i}{2}=\frac{\vartheta^{\prime}}{\vartheta}\Bigl(\frac{1}{2}F(z^{(-)}_{0},z^{(+)}_{0})-\frac{1}{4},\tau\Bigr),$
which is derived by comparing the poles $z_{0}=\pm 1,$ $\pm a_{\phi}^{1/2}$ and $\infty^{+}$ $(\in\Pi_{a_{\phi},+})$ on $\Pi_{a_{\phi}}$ (cf. [22, p.513], [28, (3.5)]). [^2] For $z_{0}=\infty^{+}$ by using $\int^{\infty}_{0}w(z)^{-1}dz=\omega_{\mathbf{b}}/2$, it is shown that $F(z_{0}^{(-)},z_{0}^{(+)})=\tau-1/2$ and the residues of the poles on both sides coincide. Similarly for each of $z_{0}=\pm 1$ and $z_{0}=\pm a_{\phi}^{1/2}$ the poles on the left-hand side are cancelled out, since $F(z^{(-)}_{0},z^{(+)}_{0})=2\sqrt{2}(1-a_{\phi})^{-1/2}(\omega_{\mathbf{a}})^{-1}t(1+o(1))$, say, for $z_{0}=1+t^{2}\to 1$. The difference of both sides is a constant, and setting $z_{0}=0$ we obtain this relation. In what follows let us adopt the convention that the path of the integral $\omega_{\mathbf{a}}F(z^{(-)}_{0},z^{(+)}_{0})=\int^{z^{(+)}_{0}}_{z^{(-)}_{0}}w(z)^{-1}dz$ passes through $a_{\phi}^{1/2}$, i.e. the left end of the cut $a^{1/2}_{\phi},1]$. Then
@@
and hence $F(-z^{(-)}_{0},-z^{(+)}_{0})=F(z^{(-)}_{0},z^{(+)}_{0})+1.$ From these relations with $_{0}=z_{+}, zz_{-}$ such that $z_{+}+z_{-}=O(t^{-1})$ (cf. (5.5)), we derive the following.
@@ 5.3. Expression of $B_{\phi}(t)$
Recall that $a_{\phi}(t)=A_{\phi}+t^{-1}B_{\phi}(t),$ $B_{\phi}(t)=O(1)$ in the domain $S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})$, where $A_{\phi}$ is a unique solution of the Boutroux equations (2.3). The quantity $B_{\phi}(t)$ is written in terms of
@@
where $w(A_{\phi},z)=\sqrt{(A_{\phi}-z^{2})(1-z^{2})},$ and $\mathbf{a},$ $\mathbf{b}$ are basic cycles on $\Pi_{A_{\phi}}$. Observing that
@@
In $S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})$, $B_{\phi}(t)$ is bounded, and
@@
in which $\hat{F}$ denotes $F$ corresponding to $\hat{\tau}=(-\omega_{\mathbf{a}})/\omega_{\mathbf{b}}.$ Since $\mathrm{Re\,}\int_{\mathbf{a},\mathbf{b}}W_{1}(z)dz$ are bounded in $S_{*}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1}),$ the Boutroux equations (2.3) with $A_{\phi}$ are equivalent to the boundedness of $\mathrm{Re\,}(e^{i\phi}\Omega_{\mathbf{a}}B_{\phi})$ and $\mathrm{Re\,}(e^{i\phi}\Omega_{\mathbf{b}}B_{\phi})$, namely, the boundedness of $B_{\phi}(t).$
@@
The derivation of the last equality is similar to that of Legendre’s relation [14, 43].
@@
By the boundedness of $B_{\phi}(t),$ $\omega_{\mathbf{a},\,\mathbf{b}}=\Omega_{\mathbf{a},\,\mathbf{b}}+O(t^{-1})$ for $\mathbf{a},$ $\mathbf{b}\subset\Pi_{a_{\phi}}\cap\Pi_{A_{\phi}}.$ From Proposition 5.5 and Corollary 5.4, it follows that
@@
Observe that the contour $z_{+}^{(-)},z_{+}^{(+)}]^{\sim}$ on $\Pi_{a_{\phi}}\cap\Pi_{A_{\phi}}$ may be decomposed into $2p_{+}(t)\mathbf{a}\cup 2q_{+}(t)\mathbf{b}\cup\mathbf{a}_{+}\cup\mathbf{a}_{-},$ where the length of $\mathbf{a}_{\pm}$ is $\ll 1$. Using (5.10) and Lemma 5.8, we have
@@
$\ll$ $|\mathcal{E}_{\mathbf{b}}j_{\mathbf{a}}-\mathcal{E}_{\mathbf{a}}j_{\mathbf{b}}|+O(t^{-1})=2\Bigl|\frac{\partial}{\partial A_{\phi}}(\mathcal{E}_{\mathbf{a}}\Omega_{\mathbf{b}}-\mathcal{E}_{\mathbf{b}}\Omega_{\mathbf{a}})\Bigr|+O(t^{-1})\ll t^{-1}$
@@
Let $y(t)$ be a function satisfying (4.3), and let $(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}$ be such that $\mathfrak{m}_{\phi}m^{0}_{21}m^{1}_{12}\not=0.$ Suppose that $0<|\phi|<\pi/2.$
@@
The same argument as in the derivation of Proposition 5.6 leads the following.
@@
Justification   The justification of (6.3) as a solution of (PV) is carried out along the argument in [28, pp. 105–106, pp. 120–121]. Let $\mathcal{M}=(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta{\infty})}$ be such that $\mathfrak{m}_{\phi}m^{0}_{21}m^{1}_{12}\not=0$ (cf. Remark 2.2). Relation (6.3) and Corollary 6.1 provide the leading terms $y_{\mathrm{as}}=y_{\mathrm{as}}(\mathcal{M},t)$ and $(B_{\phi})_{\mathrm{as}}=(B_{\phi})_{\mathrm{as}}(\mathcal{M},t)$ without $O(t^{-\delta})$. Viewing (4.3), we set $2y^{*}_{\mathrm{as}}=2y^{*}_{\mathrm{as}}(\mathcal{M},t)=\sqrt{\varphi(t,y_{\mathrm{as}},(B_{\phi})_{\mathrm{as}})}$ with
$\varphi(t,y,B)=$ $e^{2i\phi}y(4y+(1-A_{\phi})(y-1)^{2})$
@@
where the branch of the square root is chosen in such a way that the leading term of $y^{*}_{\mathrm{as}}$ is compatible with $(d/dt)y_{\mathrm{as}}$. Then $(y_{\mathrm{as}},y^{*}_{\mathrm{as}})=(y_{\mathrm{as}}(\mathcal{M},t),y^{*}_{\mathrm{as}}(\mathcal{M},t))$ satisfies (4.3) with $B_{\phi}(t)=(B_{\phi})_{\mathrm{as}}(\mathcal{M},t)$ in $\hat{S}(\phi,t_{\infty},\kappa_{0},\delta_{2})$, and the relation $\mathcal{M}=\mathcal{M}(t,y_{\mathrm{as}},y^{*}_{\mathrm{as}})$ [28, (2.28)]. Here $\hat{S}(\phi,t_{\infty},\kappa_{0},\delta_{2})=\{t\,|\,\,\mathrm{Re\,}t>t_{\infty},|\mathrm{Im\,}t|<\kappa_{0}\}\setminus\bigcup_{\sigma\in Z_{\infty}\cup Z_{0}}\{|t-e^{-i\phi}\sigma|\leq\delta_{2}\}$ with $Z_{\infty}=\{e^{i\phi}t\,|\,\,y_{\mathrm{as}}(t)=1\}=x_{0}+\Omega_{\mathbf{a}}\mathbb{Z}+2\Omega_{\mathbf{b}}(\mathbb{Z}+\tfrac{1}{2}),$ $Z_{0}=\{e^{i\phi}t\,|\,\,y_{\mathrm{as}}(t)=0,\infty\}=x_{0}+\Omega_{\mathbf{a}}(\mathbb{Z}+\tfrac{1}{2})+2\Omega_{\mathbf{b}}(\mathbb{Z}+\tfrac{1}{2}),$ and $\mathcal{M}(t,y,y^{*})$ is a collection of explicit functions $(t,y,y^{*})$ resulting from the WKB procedure. [^3] In our case $(y_{\mathrm{as}},y^{*}_{\mathrm{as}})$ may be replaced with $(y_{\mathrm{as}},(B_{\phi})_{\mathrm{as}})$, and then $\mathcal{M}(t,y_{\mathrm{as}},y^{*}_{\mathrm{as}})$ consists of, say, $e^{i\phi}\pi it-\pi i\Omega_{\mathbf{a}}(F^{*}(z^{(-)}_{\mathrm{as}},z^{(+)}_{\mathrm{as}})+\frac{1}{2})-\Omega_{\mathbf{a}}\pi i(\frac{3}{2}\theta_{\infty}+1),$ $\frac{1}{4}e^{i\phi}(t\mathcal{E}_{\mathbf{a}}+\frac{1}{2}\Omega_{\mathbf{a}}(B_{\phi})_{\mathrm{as}})-(\vartheta^{\prime}/\vartheta)(\frac{1}{2}F^{*}(z^{(-)}_{\mathrm{as}},z^{(+)}_{\mathrm{as}})-\frac{1}{4},\Omega_{\mathbf{b}}/\Omega_{\mathbf{a}})+\pi i(\theta_{\infty}+1)$, $z^{(+)}_{\mathrm{as}}=(y_{\mathrm{as}}+1)/(y_{\mathrm{as}}-1),$ which are main parts of (6.2) and of the equation of Proposition 5.6 without $O(t^{-\delta})$. The monodromy data $\mathcal{M}_{\mathrm{as}}(t)$ for system (3.4) with $\mathcal{B}(t,\lambda)$ containing $(y_{\mathrm{as}},y^{*}_{\mathrm{as}})$ is given by $\mathcal{M}_{\mathrm{as}}(t)=\mathcal{M}(t,y_{\mathrm{as}},y^{*}_{\mathrm{as}})+O(t^{-\delta})$ [28, (2.27)] as a result of the repeated WKB procedure. Thus, for $|t|\geq T_{0}$, we have $\|\mathcal{M}_{\mathrm{as}}(t)-\mathcal{M}\|\leq Ct^{-\delta}$ valid uniformly in a neighbourhood of $\mathcal{M}$, where $C$ and $T_{0}$ are independent of $\mathcal{M}$ [28, (2.26)]. Then the justification scheme [26] applies to our case (see also [12, Theorem 5.5]). This justification combined with the maximal modulus principle in each neighbourhood of $\sigma\in Z_{0}$ leads us Theorem 2.1 in $S(\rho,t_{\infty},\kappa_{0},\delta_{0}).$